### v1

# Reinforcing Question Answering Agents with Minimalist Policy Gradient Optimization

Abstract
Large Language Models (LLMs) have demonstrated remarkable versatility, due to the lack of factual knowledge, their application to Question Answering (QA) tasks remains hindered by hallucination. While Retrieval-Augmented Generation mitigates these issues by integrating external knowledge, existing approaches rely heavily on in-context learning, whose performance is constrained by the fundamental reasoning capabilities of LLMs. In this paper, we propose Mujica, a Multi-hop Joint Intelligence for Complex Question Answering, comprising a planner that decomposes questions into a directed acyclic graph of subquestions and a worker that resolves questions via retrieval and reasoning. Additionally, we introduce MyGO (Minimalist policy Gradient Optimization), a novel reinforcement learning method that replaces traditional policy gradient updates with Maximum Likelihood Estimation (MLE) by sampling trajectories from an asymptotically optimal policy. MyGO eliminates the need for gradient rescaling and reference models, ensuring stable and efficient training. Empirical results across multiple datasets demonstrate the effectiveness of Mujica-MyGO in enhancing multi-hop QA performance for various LLMs, offering a scalable and resource-efficient solution for complex QA tasks.
1 Introduction
The advent of Large Language Models (LLMs) ouyang2022training ; achiam2023gpt has redefined expectations for Artificial Intelligence. Despite showing remarkable versatility, directly using LLMs to tackle Question Answering (QA) tasks remains challenging due to their lack of factual knowledge. Without sufficient information, LLMs may hallucinate and generate plausible answers, thereby undermining the reliability of QA systems. An intuitive solution involves injecting massive up-to-date factual knowledge into LLM’s internal parameters through continued pre-training or finetuning. However, such a process not only demands substantial computational resources, but also may result in the catastrophic forgetting problem kirkpatrick2017overcoming , undermining the models’ previously acquired knowledge. To address this issue, Retrieval-Augmented Generation (RAG) guu2020retrieval ; lewis2020retrieval , a technique enabling LLMs to incorporate external knowledge via prompts, has drawn significant attention among researchers.
To effectively address complex multi-hop questions, most advanced RAG systems PoG ; Search-o1 ; RoG ; shao2023enhancing ; khattab2022demonstrate adopt a divide-and-conquer strategy. This involves decomposing the question, retrieving information for each sub-question, and synthesizing a final answer. Such a system can be naturally conceptualized as an agent taking charge of making a plan on how to answer the question, interacting with external data sources, and conducting necessary reasoning ToG ; ToG2 ; Search-o1 . Since the backbone LLMs are not specifically pre-trained on these subtasks, existing approaches primarily rely on in-context demonstration as the sole guidance. However, in-context learning (ICL) ICL  ICL has several intrinsic limitations in the QA task. First, including few-shot demonstrations requires the LLM to process more tokens for every question, which not only reduces inference speed but also increases the risk of performance degradation due to attention dispersion caused by excessively sequence lengths. Second, for different types of complex questions, one may need to provide demonstrations tailored to the specific question type to ensure effectiveness. Furthermore, in-context learning operates as a form of analogy-based reasoning, which does not improve the underlying fundamental capability of LLMs to handle QA tasks. Consequently, the model may struggle with uncommon questions in specialized domains. Therefore, to develop a truly powerful and efficient QA agent, applying post-training on LLMs becomes essential.
Training QA agents remains a challenging task. While ground truth answers are typically available, the golden reasoning paths required to derive these answers are often absent in real-world scenarios and mainstream QA datasets yang2018hotpotqa ; trivedi2021musique ; QALD-10 . Due to the lack of intermediate supervision signals, directly applying supervised finetuning (SFT) on QA agents becomes infeasible. A naive and intuitive solution is to employ larger-scale LLMs to generate reasoning paths and transfer the knowledge into smaller models through knowledge distillation. However, this approach is computationally expensive and lacks scalability, making it impractical for further improving the most advanced LLMs. Considering the aforementioned limitations, reinforcement learning (RL) emerges as a promising alternative, allowing LLMs to autonomously improve their capabilities. Recent works RAG-Gym ; IM-RAG proposes to employ the intrinsic reasoning capability of LLMs to generate “pseudo” golden reasoning paths. As training progresses, the quality of the generated reasoning paths generally improves, gradually approaching the effectiveness of knowledge distillation.
Figure 1: The end-to-end architecture of the proposed Mujica framework.
Nevertheless, directly employ existing RL solutions for LLM post-training (e.g., PPO schulman2017proximal and GRPO shao2024deepseekmath ) on complex question answering tasks is resource-intensive and inefficient. In order to stabilize the training process, these approaches schulman2017proximal ; shao2024deepseekmath ; team2025kimi ; guo2025deepseek apply gradient rescaling for individual data samples, which is primarily achieved through importance sampling and clipping mechanism. However, gradient rescaling lowers the contribution of certain useful training samples, particularly for those with higher rewards. Moreover, deploying proximity optimization-based RL solutions requires additional resources to hold a reference model, which further increases the computational complexity. It is also worth noting that the aforementioned approaches mainly focus on mathematical problems or code generation, which are typical one-turn generation scenarios. When being adapted to multi-turn complex question answering, the algorithmic and implementation inefficiencies will become more pronounced.
In this work, we approach the design of a question answering agent from a conversational perspective, and devise our solution Mujica: a Multi-hop Joint Intelligence for Complex Question Answering. Mujica consists of two modules: 1) planner, and 2) worker, each of which serve as a user and an assistant for each other. The planner extracts subquestions, organizes them into a directed acyclic graph (DAG), and iteratively calls the worker to solve each subquestion based on their interdependencies. The worker serves as a mini-RAG system responsible for interacting with external databases and answering each subquestion assigned by the planner. This design allows conditionally independent subquestions to be resolved simultaneously, thereby improving efficiency.
Furthermore, we rethink the cons of policy gradient optimization in RL, and accordingly propose Minimalist policy Gradient Optimization (MyGO) to optimize Mujica. Unlike previous methods focusing on improving the policy gradient designs such as clipping importance sampling ratios schulman2017proximal or tweaking the advantage functions, MyGO shifts the burden to the improvement of sampling strategy. By sampling trajectories from the asymptotically approximate optimal policy, we can directly use Maximum Likelihood Estimation (MLE) for policy updates, which allows MyGO to be easily implemented with standard supervised finetuning frameworks. In addition, the utilization of MLE makes MyGO stable in training, and hyper-parameter-tuning-friendly. We provide detailed theoretical justification for the validity of MyGO. Our empirical experiments on various datasets show that Mujica-MyGO effectively improves the multi-hop question answering performance for different LLMs.
2 Question Answering Agent
2.1 Problem Definition
This work addresses the task of Open-Domain Multi-Hop Question Answering (MHQA). QA systems aim to provide correct answers to given questions. Compared with single-hop QA, MHQA necessitates reasoning across and aggregating information from multiple sources (hops) to derive the final answer mavi2024multi , which renders MHQA a particularly challenging task. We focus on a more general setting, the open-domain setting, in this work, where relevant information must be retrieved from large-scale corpora (e.g., knowledge graphs, text collections). Further details are provided in the Appendix.
The inherent complexity of MHQA, characterized by multi-stage planning requirements and iterative evidence synthesis, necessitates framing the solution mechanism as an autonomous agent. We define an agent as an entity capable of planning, taking actions based on the interactions with the environment, and engaging in sequential reasoning liu2025advances . We observe that LLMs equipped with RAG perform functions analogous to such an agent. Specifically, LLMs determine information requirements and search strategies (planning), formulate search queries (acting), and synthesize final answers from gathered evidence (reasoning). Recognizing this functional correspondence, we position the LLM itself as the core agent within our MHQA system.
Notation
Let $\mathcal{Q}$ be the set of all multi-hop questions, $\mathcal{Q}^{\prime}\subset\mathcal{Q}$ be the set of all sing-hop questions, $\mathcal{M}$ be the set of all LLM agents, $\mathcal{G}$ be the set of all dependency graphs. We define the answer of question $q\in\mathcal{Q}$ as a probability distribution $a_{q}$ and actions as functions associated with agents, e.g., $\mathtt{act}(\cdot|m)$, where $\mathtt{act}$ is a placeholder for actual actions and $m\in\mathcal{M}$. When the context is clear, we omit the notation of agent, i.e., we denote simply by $\mathtt{act}(\cdot)$.
2.2 Multi-hop Joint Intelligence for Complex Question Answering (Mujica)
Figure 1 shows the overall pipeline of the proposed agentic QA framework. Given a complex question $q$, our QA agent aims to decompose it into a series of simple subquestions $\{q_{1},q_{2},\cdots,q_{n}\}$, and progressively answers each subquestion based on their internal dependencies until obtaining the final answer $a$. Let $\mathcal{I}$ denotes the instruction prompts, $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n}\}$ denote the supporting contexts of each subquestion $q_{i}$, and $a_{i}$ an answer to it. Inspired by Search-o1 , we can formulate the answer generation process as the following objective function:
${}f_{\text{QA}}(q,I,\mathcal{C})=P(a|q,I,\mathcal{C})=P(q_{1}|q,\mathcal{I})\cdot\prod_{i=2}^{n}{P(q_{i}|a_{<i},q_{<i},q,\mathcal{I})}\cdot\prod_{i=1}^{n}{P(a_{i}|q_{i},\mathcal{C}_{i},\mathcal{I})},$ (1)
where $q_{<i}$ and $a_{<i}$ denotes all preceding subquestions of subquestion $q_{i}$ and their corresponding answers. Based on the equation above, a QA agent can be naturally decoupled into two specialized modules that will be introduced in the following subsections:
  1. 1)
    Planner. Being responsible for planning how to answer a complex question, the planner determines the series of subquestions to be addressed, identifies what specific information needs to be retrieved, and adjusts the plan based on the answer of each subquestion.
  2. 2)
    Worker. Acting as a mini-RAG system, the worker directly interacts with the retriever and tackles specific sub-questions identified by the planner.
Planner and worker interact with each other through a conversational process, each serving as both user and assistant for one another. We denote our agentic QA framework as Multi-hop Joint Intelligence for Complex Question Answering (Mujica).
Planner: Planning Subquestions as a Directed Acyclic Graph.
Recent work (e.g., Search-o1 Search-o1 , PoG PoG ) has demonstrated that performing chain-of-thought reasoning through Eq. 1 is capable of deriving correct answers. However, in real-world scenarios, the subquestions posed are not naturally shaped as a sequence, since not all subquestions depend on answers to preceding ones. As shown in Fig. 2, there might be two conditionally independent subquestions, $\text{S}_{2,1}$ and $\text{S}_{2,2}$, which are dependent in the later subquestion $\text{S}_{4,1}$. The dependency relations form a directed acyclic graph (DAG). Directly applying Eq. 1 to such a reasoning process can be both inefficient and suboptimal. Therefore, to effectively handle DAG dependency graphs, we allow Mujica planner to ask subquestions in multiple iterations sequentially, where the answers of subquestions at an iterations can become dependent in later iterations. In each iteration, Mujica will simultaneously ask multiple conditionally independent subquestions.
Figure 2: QA as DAG.
More specifically, following the previous work ReAct , in each iteration, the planner will execute two steps: think and action. In the think step, at the $i$-th iteration, the planner summarizes the information obtained in the previous iterations and judges whether the supporting information it has gathered is sufficient to answer the complex question $q$. If all necessary information is available, the planner generates its conclusion for the entire complex question. Otherwise, the planner will excute the action step - it formulates the subquestions and outsources them to a worker to retrieve the corresponding information from the environment. After gathering the information from the worker, the planner will move to the next iteration, and this process continues until obtaining the final answer.
Worker: Interacting with the External Environments.
As aforementioned, the worker is responsible for handling interactions with the external environments, and answers questions assigned by the planner. Specifically, we utilize the same LLM-agent with a mini-RAG system. For each simple question $q_{i}$, the worker invokes an external retriever to fetch the top $k$ most relevant supporting contexts from a KG or a document corpus.  [^1] We employ bge-large-en-v1.5 sentence transformer as the external retriever. Then, the worker employs the LLM to review the target of the question, examine each retrieved context and accordingly anwser the subquestions outsourced by the planner. Specifically, the LLM is instructed to explicitly illustrate its thoughts, and select out relevant contexts by providing their indexes. Detailed prompts for planner and worker are provided in the Appendix.
3 Reinforcement Finetuning with Minimalist Policy Gradient Optimization
Figure 3: The proposed Minimalist Policy Gradient Optimization framework.
3.1 Preliminary: Reinforcement Learning Setting
Our QA agent is implemented within a conversational framework. A conversation (chat) is modeled as a finite sequence of interactions between the environment and the agent, denoted as $(e_{0},a_{0},e_{1},a_{1},\dots,e_{T},a_{T})$. The initial environment message $e_{0}$ comprises the user’s question and any necessary system prompts. For each turn $t\in[0,T]$, $a_{t}$ is agent’s action, and $e_{t+1}$ is the subsequent environment message. We define the state $s_{t}$ at turn $t$ as the history of all preceding interactions, $s_{t}=(e_{0},a_{0},e_{1},a_{1},\dots,a_{t-1},e_{t})$. Then, a trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},\dots,s_{T},a_{T})$ to the final answer of the question can be denoted as the sequence of states and actions in each step. This formulation adheres to a Markov Decision Process (MDP). Since the golden reasoning path to the ground truth answer is not available in real-world scenarios and most datasets, we do not have supervision signals for each action $a_{t}$ to conduct supervised finetuning. The only supervision signal available to estimate the reward $r(\tau)\in\mathbb{R}$ for the entire trajectory $\tau$ is the final answer, which naturally corresponds to terminal reward in an RL setting.
We consider a policy-based method - the QA agent adopts a policy $\pi_{\theta}$ to generate reasoning trajectories, where the parameters $\theta$ are optimized using RL techniques. The objective of the RL is to find the best policy that maximizes the expected cumulative reward:
$\mathcal{J}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[r(\tau)]\;,$ (2)
If trajectories sampled from $\pi_{\theta}$ are differentiable w.r.t. $\theta$, the gradient of the optimization objective can be estimated using the Vanilla Policy Gradient (VPO) (williams1992simple, ):
$\nabla_{\theta}\mathcal{J}(\theta)=\underset{\tau\sim\pi_{\theta}}{\mathbb{E}}\left[r(\tau)\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right]\;.$ (3)
3.2 Reinforcement Learning for Large Language Models
Even though the VPO (Eq. 3) estimator is unbiased, it often exhibits high gradient variance, which potentially renders the training process unstable. To address this issue, many previous approaches schulman2015trust ; schulman2017proximal propose to replace the raw trajectory reward $r(\tau)$ with the Generalized Advantage Estimation (GAE) (schulman2016high, ). These RL techniques, particularly Proximal Policy Optimization (PPO) schulman2017proximal and its variants, have been successfully applied in the post-training process for LLM (ouyang2022training, ; rafailov2023direct, ). Recently, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) has been introduced as a simplified alternative to PPO, reducing the training costs while maintaining comparable performance. The objective function of PPO and GRPO, as adapted for our QA context, can be written as follows:
$\mathcal{J}(\theta)=\underset{\tau\sim\pi_{\theta_{k}}}{\mathbb{E}}\left[A(\tau)\sum_{t=0}^{T}\min\left(\mathbf{r}_{t},\;\text{clip}\left(\mathbf{r}_{t},1-\epsilon,1+\epsilon\right)\right)\right]-\beta\mathds{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\;,$ (4)
where $\epsilon,\beta$ are hyperparameters; $\pi_{\text{ref}},\pi_{\theta_{k}}$ denote the reference policy and the behavior policy, respectively; $\mathbf{r}_{t}=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{k}}(a_{t}|s_{t})}$ is the importance sampling rescaling factor; and $A(\tau)$ is the advantage of applying trajectory $\tau$. Specifically, in PPO, the advantage function $A(\tau)$ is estimated by $r(\tau)-V(\tau)$, relying on a learnable value function $V(\tau)$. In GRPO, $A(\tau)$ is instead estimated by $\frac{r(\tau)-\bar{r}}{\text{std}(r)}$, where $\bar{r}$ and $\text{std}(r)$ are the mean and standard deviation of $r(\tau)$ for each $\tau$ obtained from the same sampling group, respectively.
3.3 Minimalist Policy Gradient Optimization (MyGO)
Despite its wide adoption, a PPO-style objective has two shortcomings, hyperparameter sensitivity and training inefficiency. The $\epsilon$ hyperparameter is highly sensitive - a small $\epsilon$ may ignore most available trajectories during training, leading to training inefficiency; while a large $\epsilon$ might lose control over the importance sampling scale, resulting in a less stable training process. The advantage estimation renormalizes the original rewards relative to either the value functions or the group average, which weakens the supervision signals from trajectories with large rewards that are helpful for training liu2024decade . It is worth mentioning that training inefficiency and instability are significant challenges when training LLMs, which is both computational expensive and time-consuming. On the other hand, compared with previous RL scenarios in robot control and video games openai2016gym , LLM setting poses new characteristic: fast simulation, deterministic environment, and trajectory-level reward li2024remax . In this scenario, sampling desired trajectories from the environment is relatively more efficient.
To this end, we propose our new RL algorithm, Minimalist Policy Gradient Optimization (MyGO), which shifts the burden from model training to trajectory sampling, in which we can fully leverage the fast and relatively inexpensive simulation of LLMs kwon2023efficient ; zheng2024sglang . Unlike previous methods, MyGO disentangles the RL training process into two phases: sampling and training. During sampling, MyGO directly samples trajectories from the asymptotically optimal distribution and consequently allows the use of Maximum Likelihood Estimation (MLE) to optimize the policy function. Notably, optimizing a model via MLE is simple, stable, and well established as a regular training technique widley used in unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training .
Sampling
We first characterize the target optimal distribution $\pi^{*}$. Alternative to the original objective $\mathcal{J}$, we consider the expected reward with entropy regularization $\mathcal{J}^{\prime}=\mathbb{E}[r(\tau)]+\alpha\mathbf{H}(\pi)$ with $\alpha\in\mathbb{R}^{+}$, which, as a Boltzmann distribution ziebart2010modeling , has a closed-form for the corresponding optimal policy, denoted as ${\pi^{\prime}}^{*}$:
[…]
where $\alpha$ is the temperature and $Z(\alpha)$ is the partition function that normalizes the distribution over the space of all possible trajectories $\mathcal{T}$. With a sufficiently small $\alpha\to 0$, we have $\mathcal{J}^{\prime}\to\mathcal{J}$, thus ${\pi^{\prime}}^{*}\to\pi^{*}$. As $\alpha\to 0$, ${\pi^{\prime}}^{*}$ becomes sharp and highly skewed, concentrating its probability mass on trajectories yielding high rewards. Therefore, given a sufficiently large threshold $K$ close to $\sup_{\tau\in\mathcal{T}}r(\tau)$, sampling trajectories $\tau$ with $r(\tau)>K$ is asymptotically equivalent to sampling from the ${\pi^{\prime}}^{*}$ and thus effectively approximating to sampling from the optimal policy ${\pi}^{*}$ of $\mathcal{J}$.
In practice, we relax the choice of $K$ by setting an update role of $K:=\max(K^{\prime},(\bar{r}_{n}/r_{\text{sup}}+1)\cdot r_{\text{sup}})$, where $r_{\text{sup}}$ denotes $\sup_{\tau\in\mathcal{T}}r(\tau)$, $\bar{r}_{n}$ is the empirical average reward of a batch of $n$ sampled trajectories, and $K^{\prime}$ is the previous threshold value which is initialized as a moderate value depending on the choice of dataset. As the policy $\pi_{\theta}$ improved, the $K$ will be asymptotically close to $\sup_{\tau\in\mathcal{T}}r(\tau)$, thus making the sampling distribution close to ${\pi^{\prime}}^{*}$.
We highlight that this approach is practically feasible for the reward landscapes for tasks of LLM agent, which differ significantly from that in complex robotic control problems. For LLM agents, trajectories frequently culminate in outcomes that are clearly classifiable as successful or unsuccessful. Empirically, we observe that well-prompted LLMs can readily generate a significant proportion of such successful trajectories, which makes directly sampling from $\pi^{*}$ feasible.
Let $\pi^{>K}$ denote the distribution that consists of trajectories with $r>K$. To clarify, we provide two propositions as follows, where the proofs can be found in the Appendix:
Proposition 3.1.
Given $\delta\in\mathbb{R}$, $\alpha\in R$, there exists $K\in\mathbb{R}^{+}$ such that the following inequality holds:
$\mathds{D}_{\text{KL}}(\pi^{>K}||\pi^{*})<\delta\;,$
Proposition 3.2.
$\mathrm{Var}(\pi^{>K})\propto 1/Z(\alpha)^{2}\;.$
Proposition 3.1 states that if $K$ is sufficiently large, that the $\pi^{>K}$ can always approximate $\pi^{*}$. Proposition 3.2 says the variance of $\pi^{>K}$ is small, considering that $Z$ is usually a large value. These two propositions guarantee the proximity and stability of our proposed sampling strategy.
Learning
Once a sufficiently large dataset of optimal trajectories $\mathcal{D}$ has been collected, the optimization of policy $\pi_{\theta}$ can be simplified to maximizing the log-likelihood $\mathcal{L}$ of these trajectories. This yields the following MLE gradient:
$\nabla_{\theta}\mathcal{L}$ $=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\nabla_{\theta}\log\pi_{\theta}(\tau_{i})=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\sum_{j=0}^{|\tau_{i}|}\nabla_{\theta}\mathds{1}(o_{j}^{i})\log\pi_{\theta}(o_{j}^{i}|o_{<j}^{i})\;,$ (6)
where $o_{j}^{i}$ denotes the $j$-th token from $\tau_{i}$ and $o_{<j}^{i}$ denotes the sequence of tokens preceding $o_{j}$, $\mathds{1}(o_{j})$ is a characteristic function indicating whether $o_{j}$ is a token from the agent. The last equality holds because the action likelihood in language setting could be further elaborated with token likelihood, i.e, $\log\pi_{\theta}(a_{t}|s_{t})=\sum_{j}\log\pi_{\theta}(o_{j}|o_{<j})$.
This MLE objective, Eq. 6, is equivalent to a standard cross-entropy loss function, widely used in both unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training of language models. Compared to PPO-style algorithms, Eq. 4, MLE does not incorporate an advantage function, policy ratio clipping, and importance sampling. Similarly, unlike methods such as VPO, Eq. 3, MLE directly uses the curated data $\mathcal{D}$ for training and does not explicitly use reward values in the loss computation. We empirically observed that MLE leads to more stable training. We explain this increased stability by several factors: (i) the dataset $\mathcal{D}$ consists of trajectories sampled from a more stationary (approximated optimal) distribution, reducing variance compared to on-policy RL where the data distribution shifts rapidly; (ii) the direct use of a log-likelihood objective (cross-entropy) is often well-behaved and less sensitive to reward scaling issues that can affect RL algorithms; and (iii) the avoidance of importance sampling ratios, which can introduce instability if they become too large or small. A detailed theoretical justification for these empirical observations is left for future work.
Compared with previous works, like RAFT dongraft and STaR zelikman2022star , who try to simplify RL in LLM post-training, we use a progressive threshold to select data. More importantly, we provide a theoretical justification that elucidates why this approach, yet relatively simple, yields strong performance, offering novel insights into its efficacy.
4 Experiments
4.1 Environment Design
Datasets and Evaluation Metrics
To evaluate the effectiveness of our proposed agent workflow (Mujica) and training method (MyGO) on complex question answering tasks, we utilize four representative benchmark datasets: 2Wiki-MultihopQA (2Wiki) (ho2020constructing, ), QALD-10 (QALD-10, ), HotpotQA (yang2018hotpotqa, ), and MuSiQue (trivedi2021musique, ).
2Wiki and QALD-10 are KBQA benchmarks; HotpotQA and MuSiQue are text-based QA dataset. We use 2Wiki and HotpotQA for training, QALD-10 and MuSiQue for zero-shot evaluation. We assess QA performance using Exact Match (EM) and F1 score. For each dataset, these metrics are computed using the official evaluation scripts provided by the respective dataset creators. Comprehensive dataset statistics and precise definitions of the evaluation metrics are detailed in Appendix.
Environment Setting
To evaluate our agent’s capabilities across different data modalities, we derive two experimental environments from the 2Wik dataset: 2Wiki-KG and 2Wiki-Text. This setup allows us to agent’s performance on both structured (KGs) and unstructured (text) data. We use the F1 score as reward function for all environments.
The 2Wiki-KG environment is centered around the Wikidata KG (Wikidata, ). A KG is composed of factual statements represented as triples, such as (Donald Trump, position held, President of the United States). When the agent visits an entity, it can access all triples for which that entity is the head. The agent’s task involves traversing the KG by interpreting the semantic relationships encoded in these triples. For example, to answer the question "Who is the mother of the director of film Polish-Russian War (Film)?", the agent must identify and follow the correct reasoning path, such as: Polish-Russian War, director, Xawery Żuławski, mother, Małgorzata Braunek. To emphasize the reasoning aspect, we follow prior work (ToG, ; ToG2, ) by providing the agent with the gold topic entities extracted from the dataset as the entry.
We establish the 2Wiki-Text environment to further evaluate our agent’s ability to interact with text-based search engines. The original 2Wiki dataset provides ten supporting passages for each question, which contain the necessary information to derive the answer and the "distractor setting" is designed to test multi-hop reasoning abilities without retrieval. In our 2Wiki-Text environment, we adapt this by aggregating all passages to form a large corpus. This approach, following prior work (IRCoT, ; aguerrebere2023similarity, ), simulates an open-domain retrieval task. The agent’s objective is to answer questions by issuing search queries and then synthesizing information from the retrieved passages.
Beyond the 2Wiki environments, we extend our text-based evaluation using the HotpotQA dataset, from which we construct two additional settings: Hotpot and Hotpot-Kimi. In Hotpot, we folow the vanilla setting that uses the Wikipedia dump provided by the dataset yang2018hotpotqa , which only includes the title and the introductory paragraph. However, we found that the retriever karpukhin2020dense is the bottleneck for QA performance in the HotpotQA dataset. To isolate and more directly assess the agent’s reasoning capabilities independent of retrieval imperfections, we introduce the Hotpot-Kimi environment. Similar to but different from the distractor settings (yang2018hotpotqa, ; ho2020constructing, ), in the Hotpot-Kimi environment,the agent is directly provided with the ten gold supporting passages associated with each question, thereby guaranteeing access to all necessary evidence. However, its primary purpose here is to establish an idealized retrieval condition. Therefore, the gold passages are not directly fed into the QA-agent like distractor setting. Detailed examples and prompts for each environment are provided in Appendix.
4.2 Main Results
Table 1: Comparative Performance of Models on 2Wiki-KG and 2Wiki-Text Environments.
Type Model 2Wiki-KG 2Wiki-Text
EM F1 EM F1
Few-Shot GPT-4.1 78.50 84.24 30.70 52.14
Few-Shot Qwen2.5-7B 57.49 62.31 12.41 19.71
WarmUp 74.93 80.74 50.18 56.25
MyGO 77.63 84.15 53.17 59.62
Few-Shot Llama3.1-8B 77.70 82.09 37.83 46.53
WarmUp 82.00 85.90 58.03 64.81
MyGO 85.93 91.61 58.88 65.88
We evaluate our QA agent under three distinct settings: Few-shot, WarmUp, and MyGO. In the Few-shot setting, the agent is provided with a small number of illustrative examples via in-context learning within the prompt. Following common practice, we then perform a WarmUp phase. This phase is to adapt the base model beyond reliance on few-shot exemplars. For this purpose, we fine-tune the model on 1k samples with EM of 1. The MyGO training then commences, using this warmed-up model as its initialization. We experiment with two open-source models: Qwen2.5-7B-Instruct (Qwen) and Llama-3.1-8B-Instruct (Llama).
Table 1 presents the performance of these models on the 2Wiki dataset. Agents trained with MyGO consistently demonstrate strong performance. In the 2Wiki-KG environment, MyGO-Qwen achieves performance on par with a few-shot prompted GPT-4.1, while the Llama model significantly surpasses this baseline. In the 2Wiki-Text environment, both MyGO-trained Qwen and Llama models substantially outperform the GPT-4.1 few-shot baseline. Despite the same underlying questions, overall performance in the 2Wiki-Text environment is notably lower than in 2Wiki-KG. This disparity can be attributed to the differing data modalities and interaction paradigms. Interacting with KG may more closely resemble navigating discrete choices (e.g., selecting entities or relations), potentially reducing the burden on complex retrieval or synthesis from noisy search results. Conversely, text-based interaction requires robust open-retrieval over the corpus where explicit relationships between retrieved passages are often absent, demanding more sophisticated synthesis capabilities. Nevertheless, the consistent improvements observed affirm MyGO’s efficacy across these varied settings.
Table 2: Comparative Performance of Qwen on Hotpot and Hotpot-Kimi Environment.
Type Model Hotpot Hotpot-Kimi
EM F1 EM F1
Few-Shot GPT-4.1 29.80 48.09 33.50 53.35
Few-Shot Qwen2.5-7B 15.77 27.96 15.60 32.70
WarmUp 40.55 52.35 52.51 66.04
MyGO 41.54 53.79 54.07 68.48
Table 3 details the results on the HotpotQA dataset. Notably, even the WarmUp models exhibit strong performance, outperforming the few-shot GPT-4.1 baseline. However, the performance gain from MyGO over the WarmUp models is less pronounced in the Hotpot environment. We attribute this observation to two limitations inherent in this setup. Firstly, the efficacy of the dense retriever can be a bottleneck; essential evidence may not be surfaced for some queries. Secondly, the HotpotQA evaluation protocol has no answer aliases. This is problematic because LLMs frequently generate answers that are semantically equivalent to the gold answer but lexically different, which are penalized by strict string-matching metrics. To mitigate the impact of imperfect retrieval, we further evaluate using the Hotpot-Kimi environment, which is designed to simulate near-perfect retrieval by providing gold supporting passages. In this idealized Hotpot-Kimi setting, MyGO demonstrates substantial and consistent performance improvements over the WarmUp models, underscoring its effectiveness when reasoning capabilities can be more directly assessed. A more comprehensive comparison is provided in Appendix.
4.3 Out-Of-Distribution Evaluation
[…]
Type Model H $\to$ M 2W $\to$ Q
EM F1 EM F1
Few-Shot Qwen 2.5-7B 5.34 17.81 27.92 37.11
MyGO 26.11 35.92 39.85 49.73
To assess generalization capabilities, we conduct out-of-distribution (OOD) evaluations where models are RL post-trained on one dataset and evaluated on another one. Specifically, we examine cross-dataset transfer from HotpotQA to Musique (text-based), and from 2Wiki-KG to Qald-10 (KG-based). As shown in Table 3, MyGO demonstrates strong OOD performance. As presented in Table 3, models trained with MyGO demonstrate robust OOD generalization. Specifically, applying MyGO on the source datasets leads to significant performance improvements on the respective target OOD datasets when compared to the few-shot baseline. These findings indicate that MyGO does not overfit to the training data; rather, it appears to foster the acquisition of more generalizable reasoning and decision-making strategies. This capacity for effective generalization to unseen datasets aligns with a primary motivation for employing RL in the development of intelligent agents.
5 Related Works
MHQA with textual corpus.
Retrieval-based methods dominate the landscape of textual MHQA since supporting contexts from external corpora can effectively compensate for deficiencies in the inherent knowledge of language models and mitigate hallucinations. Traditional approaches primarily utilized sentence transformers bgem3 for dense retrieval or BM25 BM25 for keyword matching. Recently, extensive research has focused on optimizing the QA pipeline, particularly through the decomposition of complex questions IM-RAG ; ReAct , query rewriting and refinement RQ-RAG ; RaFe , and the iterative retrieval of relevant evidences in multiple steps SelfRAG ; ReSP ; Iter-RetRAG ; IRCoT .
Knowledge-based question answering.
Early approaches treat KBQA as a semantic parsing task, which generates and executes SPARQL queries lan-jiang-2020-query ; SPARQL2021 ; KD-CoT . These methods usually suffer from syntax errors and inaccurate entity labels, resulting in limited accuracy. Recently, researchers have shifted towards leveraging RAG and the advanced semantic understanding capabilities of LLMs to generate answers RRA . LLM-based KBQA approaches can be generally categorized into two types: iterative exploration and subgraph reasoning. Iterative exploration methods ToG ; ToG2 ; ToG-I ; KG-CoT ; PoG ; FiDeLiS ; GoG ; KARPA ; Paths-over-Graph ; BeamRetrieval ; DoG progressively retrieve and verify KG elements through multi-step reasoning, balancing interpretability with computational overhead. These approaches rely heavily on LLMs’ inherent reasoning capabilities due to the lack of supervised intermediate steps in most datasets. In contrast, Subgraph reasoning methods RoG ; GNN-RAG ; SubgraphRAG ; READS ; GCR extract relevant KG substructures in a single retrieval step for LLM-based answering, enabling efficient fine-tuning but suffering from the incompleteness of retrieved subgraphs and underlying knowledge bases.
6 Conclusion
In this work, we proposed Mujica, an interactive conversational framework for multi-hop question answering, and introduced MyGO, a novel RL optimization method that overcomes the inefficiencies of Proximity Optimization-based approaches. Mujica leverages a planner-worker structure to efficiently decompose and resolve complex questions, while MyGO improves the training efficiency of proximity optimization methods by shifting the burden from clipping importance ratio and tweaking advantage functions to trajectory sampling. Our experiments demonstrate that the proposed framework not only achieves significant performance improvements across multiple QA datasets but also exhibits strong inductive reasoning capabilities, highlighting its scalability and potential to generalize effectively to real-world scenarios.