### v2

# Reinforcing Question Answering Agents with Minimalist Policy Gradient Optimization

Abstract
Large Language Models (LLMs) show remarkable versatility but face challenges in Question Answering (QA) due to knowledge cutoffs and hallucination. While Retrieval-Augmented Generation (RAG) helps by integrating external knowledge, current methods often depend on in-context learning, which can be limited by backbone models and prompt engineering. To overcome these limitations, we introduce Mujica (Multi-hop Joint Intelligence for Complex Question Answering), a novel multi-agent collaborative framework employing a divide-and-conquer strategy for QA. This framework is trained using MyGO (Minimalist policy Gradient Optimization), a new reinforcement learning method that directly samples and learns from the optimal policy. Empirical results across multiple datasets demonstrate that the Mujica-MyGO approach significantly enhances multi-hop QA performance for various LLMs.
1 Introduction
The advent of Large Language Models (LLMs) ouyang2022training ; achiam2023gpt has redefined expectations for Artificial Intelligence, offering new opportunities for Question Answering (QA) tasks due to their impressive language understanding and generation capabilities. However, directly applying LLMs to QA is challenging because of hallucination huang2025survey and knowledge cutoffs cheng2024dated . While injecting knowledge through continue learning is a possible solution, it is computationally expensive and risks catastrophic forgetting kirkpatrick2017overcoming , where models lose previously acquired knowledge. To address this issue, Retrieval-Augmented Generation (RAG) guu2020retrieval ; lewis2020retrieval , a technique enabling LLMs to incorporate external knowledge via prompts, has drawn significant attention among researchers.
Compared with single-hop questions which have one-turn retrieve-then-answer pipeline, most advanced RAG systems PoG ; Search-o1 ; RoG ; shao2023enhancing ; khattab2022demonstrate adopt a multi-turn pipeline to effectively address multi-hop questions. This involves decomposing the question, retrieving information for each sub-question, and synthesizing a final answer. Such a system can be naturally conceptualized as an agent taking charge of making a plan on how to answer the question, interacting with external data sources, and conducting necessary reasoning ToG ; ToG2 ; Search-o1 . However, existing approaches primarily rely on in-context demonstration as the sole guidance. However, in-context learning (ICL) ICL has several intrinsic limitations in the QA task. First, including few-shot demonstrations requires the LLM to process more tokens for every question, which not only reduces inference speed but also increases the risk of performance degradation due to attention dispersion caused by excessively sequence lengths. Second, for different types of complex questions, one may need to provide demonstrations tailored to the specific question type to ensure effectiveness. Furthermore, in-context learning operates as a form of analogy-based reasoning, which does not improve the underlying fundamental capability of LLMs to handle QA tasks. Consequently, the model may struggle with uncommon questions in specialized domains. Therefore, to develop a truly powerful and efficient QA agent, applying post-training on LLMs becomes essential.
Training QA agents remains a challenging task. While ground truth answers are typically available, the golden reasoning paths required to derive these answers are often absent in real-world scenarios and mainstream QA datasets yang2018hotpotqa ; trivedi2021musique ; QALD-10 . Due to the lack of intermediate supervision signals, directly applying supervised finetuning (SFT) on QA agents becomes infeasible. A naive and intuitive solution is to employ larger-scale LLMs to generate reasoning paths and transfer the knowledge into smaller models through knowledge distillation. However, this approach is computationally expensive and lacks scalability, making it impractical for further improving the most advanced LLMs. Considering the aforementioned limitations, reinforcement learning (RL) emerges as a promising alternative, allowing LLMs to autonomously improve their capabilities. Recent works RAG-Gym ; IM-RAG proposes to employ the intrinsic reasoning capability of LLMs to generate “pseudo” golden reasoning paths. As training progresses, the quality of the generated reasoning paths generally improves, gradually approaching the effectiveness of knowledge distillation.
Figure 1: The end-to-end architecture of the proposed Mujica framework.
Nevertheless, directly employ existing RL solutions for LLM post-training (e.g., PPO schulman2017proximal and GRPO shao2024deepseekmath ) on complex question answering tasks is resource-intensive and inefficient. In order to stabilize the training process, these approaches schulman2017proximal ; shao2024deepseekmath ; team2025kimi ; guo2025deepseek apply gradient rescaling for individual data samples, which is primarily achieved through importance sampling and clipping mechanism. However, gradient rescaling lowers the contribution of certain useful training samples, particularly for those with higher rewards. Moreover, deploying proximity optimization-based RL solutions requires additional resources to hold a reference model, which further increases the computational complexity. It is also worth noting that the aforementioned approaches mainly focus on mathematical problems or code generation, which are typical one-turn generation scenarios. When being adapted to multi-turn complex question answering, the algorithmic and implementation inefficiencies will become more pronounced.
In this work, we approach the design of a question answering agent from a conversational perspective, and devise our solution Mujica: a Multi-hop Joint Intelligence for Complex Question Answering. Mujica consists of two modules: 1) planner, and 2) worker, each of which serve as a user and an assistant for each other. The planner extracts subquestions, organizes them into a directed acyclic graph (DAG), and iteratively calls the worker to solve each subquestion based on their interdependencies. The worker serves as a mini-RAG system responsible for interacting with external databases and answering each subquestion assigned by the planner. This design allows conditionally independent subquestions to be resolved simultaneously, thereby improving efficiency.
Furthermore, we rethink the cons of policy gradient optimization in RL, and accordingly propose Minimalist policy Gradient Optimization (MyGO) to optimize Mujica. Unlike previous methods focusing on improving the policy gradient designs such as clipping importance sampling ratios schulman2017proximal or tweaking the advantage functions, MyGO shifts the burden to the improvement of sampling strategy. By sampling trajectories from the asymptotically approximate optimal policy, we can directly use Maximum Likelihood Estimation (MLE) for policy updates, which allows MyGO to be easily implemented with standard supervised finetuning frameworks. In addition, the utilization of MLE makes MyGO stable in training, and hyper-parameter-tuning-friendly. We provide detailed theoretical justification for the validity of MyGO. Our empirical experiments on various datasets show that Mujica-MyGO effectively improves the multi-hop question answering performance for different LLMs.
2 Question Answering Agent
2.1 Problem Definition
This work addresses the task of Open-Domain Multi-Hop Question Answering (MHQA). QA systems aim to provide correct answers to given questions. Compared with single-hop QA, MHQA necessitates reasoning across multiple sources (hops) mavi2024multi , which renders MHQA a more challenging task. We focus on the open-domain setting, where relevant information must be retrieved from large-scale corpora (knowledge graphs or text collections). Details are provided in the Appendix.
The inherent complexity of MHQA, characterized by multi-stage planning requirements and iterative evidence synthesis, necessitates framing the solution mechanism as an autonomous agent. We define an agent as an entity capable of planning, taking actions based on the environment, and engaging in sequential reasoning liu2025advances . We observe that LLMs equipped with RAG perform functions analogous to such an agent. Specifically, LLMs determine information requirements and search strategies (planning), formulate search queries (acting), and synthesize final answers from gathered evidence (reasoning). Recognizing this functional correspondence, we position the LLM itself as the core agent within our MHQA system.
Notation
Let $\mathcal{Q}$ be the set of all multi-hop questions, $\mathcal{Q}^{\prime}\subset\mathcal{Q}$ be the set of all sing-hop questions, $\mathcal{M}$ be the set of all LLM agents, $\mathcal{G}$ be the set of all dependency graphs. We define the answer of question $q\in\mathcal{Q}$ as a probability distribution $a_{q}$ and actions as functions associated with agents, e.g., $\mathtt{act}(\cdot|m)$, where $\mathtt{act}$ is a placeholder for actual actions and $m\in\mathcal{M}$. When the context is clear, we omit the notation of agent, i.e., we denote simply by $\mathtt{act}(\cdot)$.
2.2 Multi-hop Joint Intelligence for Complex Question Answering (Mujica)
Figure 1 shows the overall pipeline of the proposed agentic QA framework. Given a complex question $q$, our QA agent aims to decompose it into a series of simple subquestions $\{q_{1},q_{2},\cdots,q_{n}\}$, and progressively answers each subquestion based on their internal dependencies until obtaining the final answer $a$. Let $\mathcal{I}$ denotes the instruction prompts, $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n}\}$ denote the supporting contexts of each subquestion $q_{i}$, and $a_{i}$ an answer to it. Inspired by Search-o1 , we can formulate the answer generation process as the following objective function:
${}f_{\text{QA}}(q,I,\mathcal{C})=P(a|q,I,\mathcal{C})=P(q_{1}|q,\mathcal{I})\cdot\prod_{i=2}^{n}{P(q_{i}|a_{<i},q_{<i},q,\mathcal{I})}\cdot\prod_{i=1}^{n}{P(a_{i}|q_{i},\mathcal{C}_{i},\mathcal{I})},$ (1)
where $q_{<i}$ and $a_{<i}$ denotes all preceding subquestions of subquestion $q_{i}$ and their corresponding answers. Based on the equation above, a QA agent can be naturally decoupled into two specialized modules that will be introduced in the following subsections:
  1. 1)
    Planner. Being responsible for planning how to answer a complex question, the planner determines the series of subquestions to be addressed, identifies what specific information needs to be retrieved, and adjusts the plan based on the answer of each subquestion.
  2. 2)
    Worker. Acting as a mini-RAG system, the worker directly interacts with the retriever and tackles specific sub-questions identified by the planner.
Planner and worker interact with each other through a conversational process, each serving as both user and assistant for one another. We denote our agentic QA framework as Multi-hop Joint Intelligence for Complex Question Answering (Mujica).
Planner: Planning Subquestions as a Directed Acyclic Graph.
Recent work (e.g., Search-o1 Search-o1 , PoG PoG ) has demonstrated that performing chain-of-thought reasoning through Eq. 1 is capable of deriving correct answers. However, in real-world scenarios, the subquestions posed are not naturally shaped as a sequence, since not all subquestions depend on answers to preceding ones. As shown in Fig. 2, there might be two conditionally independent subquestions, $\text{S}_{2,1}$ and $\text{S}_{2,2}$, which are dependent in the later subquestion $\text{S}_{4,1}$. The dependency relations form a directed acyclic graph (DAG). Directly applying Eq. 1 to such a reasoning process can be both inefficient and suboptimal. Therefore, to effectively handle DAG dependency graphs, we allow Mujica planner to ask subquestions in multiple iterations sequentially, where the answers of subquestions at an iterations can become dependent in later iterations. In each iteration, Mujica will simultaneously ask multiple conditionally independent subquestions.
Figure 2: QA as DAG.
More specifically, following the previous work ReAct , in each iteration, the planner will execute two steps: think and action. In the think step, at the $i$-th iteration, the planner summarizes the information obtained in the previous iterations and judges whether the supporting information it has gathered is sufficient to answer the complex question $q$. If all necessary information is available, the planner generates its conclusion for the entire complex question. Otherwise, the planner will excute the action step - it formulates the subquestions and outsources them to a worker to retrieve the corresponding information from the environment. After gathering the information from the worker, the planner will move to the next iteration, and this process continues until obtaining the final answer.
Worker: Interacting with the External Environments.
As aforementioned, the worker is responsible for handling interactions with the external environments, and answers questions assigned by the planner. Specifically, we utilize the same LLM-agent with a mini-RAG system. For each simple question $q_{i}$, the worker invokes an external retriever to fetch the top $k$ most relevant supporting contexts from a KG or a document corpus.  [^1] We employ bge-large-en-v1.5 sentence transformer as the external retriever. Then, the worker employs the LLM to review the target of the question, examine each retrieved context and accordingly anwser the subquestions outsourced by the planner. Specifically, the LLM is instructed to explicitly illustrate its thoughts, and select out relevant contexts by providing their indexes. Detailed prompts for planner and worker are provided in the Appendix.
3 Reinforcement Finetuning with Minimalist Policy Gradient Optimization
Figure 3: The proposed Minimalist Policy Gradient Optimization framework.
3.1 Preliminary: Reinforcement Learning Setting
Our QA agent is implemented within a conversational framework. A conversation (chat) is modeled as a finite sequence of interactions between the environment and the agent, denoted as $(e_{0},a_{0},e_{1},a_{1},\dots,e_{T},a_{T})$. The initial environment message $e_{0}$ comprises the user’s question and any necessary system prompts. For each turn $t\in[0,T]$, $a_{t}$ is agent’s action, and $e_{t+1}$ is the subsequent environment message. We define the state $s_{t}$ at turn $t$ as the history of all preceding interactions, $s_{t}=(e_{0},a_{0},e_{1},a_{1},\dots,a_{t-1},e_{t})$. Then, a trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},\dots,s_{T},a_{T})$ to the final answer of the question can be denoted as the sequence of states and actions in each step. This formulation adheres to a Markov Decision Process (MDP). Since the golden reasoning path to the ground truth answer is not available in real-world scenarios and most datasets, we do not have supervision signals for each action $a_{t}$ to conduct supervised finetuning. The only supervision signal available to estimate the reward $r(\tau)\in\mathbb{R}$ for the entire trajectory $\tau$ is the final answer, which naturally corresponds to terminal reward in an RL setting.
We consider a policy-based method - the QA agent adopts a policy $\pi_{\theta}$ to generate reasoning trajectories, where the parameters $\theta$ are optimized using RL techniques. The objective of the RL is to find the best policy that maximizes the expected cumulative reward:
$\mathcal{J}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[r(\tau)]\;,$ (2)
If trajectories sampled from $\pi_{\theta}$ are differentiable w.r.t. $\theta$, the gradient of the optimization objective can be estimated using the Vanilla Policy Gradient (VPO) (williams1992simple, ):
$\nabla_{\theta}\mathcal{J}(\theta)=\underset{\tau\sim\pi_{\theta}}{\mathbb{E}}\left[r(\tau)\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right]\;.$ (3)
3.2 Reinforcement Learning for Large Language Models
Even though the VPO (Eq. 3) estimator is unbiased, it often exhibits high gradient variance, which potentially renders the training process unstable. To address this issue, many previous approaches schulman2015trust ; schulman2017proximal propose to replace the raw trajectory reward $r(\tau)$ with the Generalized Advantage Estimation (GAE) (schulman2016high, ). These RL techniques, particularly Proximal Policy Optimization (PPO) schulman2017proximal and its variants, have been successfully applied in the post-training process for LLM (ouyang2022training, ; rafailov2023direct, ). Recently, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) has been introduced as a simplified alternative to PPO, reducing the training costs while maintaining comparable performance. The objective function of PPO and GRPO, as adapted for our QA context, can be written as follows:
$\mathcal{J}(\theta)=\underset{\tau\sim\pi_{\theta_{k}}}{\mathbb{E}}\left[A(\tau)\sum_{t=0}^{T}\min\left(\mathbf{r}_{t},\;\text{clip}\left(\mathbf{r}_{t},1-\epsilon,1+\epsilon\right)\right)\right]-\beta\mathds{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\;,$ (4)
where $\epsilon,\beta$ are hyperparameters; $\pi_{\text{ref}},\pi_{\theta_{k}}$ denote the reference policy and the behavior policy, respectively; $\mathbf{r}_{t}=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{k}}(a_{t}|s_{t})}$ is the importance sampling rescaling factor; and $A(\tau)$ is the advantage of applying trajectory $\tau$. Specifically, in PPO, the advantage function $A(\tau)$ is estimated by $r(\tau)-V(\tau)$, relying on a learnable value function $V(\tau)$. In GRPO, $A(\tau)$ is instead estimated by $\frac{r(\tau)-\bar{r}}{\text{std}(r)}$, where $\bar{r}$ and $\text{std}(r)$ are the mean and standard deviation of $r(\tau)$ for each $\tau$ obtained from the same sampling group, respectively.
3.3 Minimalist Policy Gradient Optimization (MyGO)
Despite its wide adoption, a PPO-style objective has two shortcomings, hyperparameter sensitivity and training inefficiency. The $\epsilon$ hyperparameter is highly sensitive - a small $\epsilon$ may ignore most available trajectories during training, leading to training inefficiency; while a large $\epsilon$ might lose control over the importance sampling scale, resulting in a less stable training process. The advantage estimation renormalizes the original rewards relative to either the value functions or the group average, which weakens the supervision signals from trajectories with large rewards that are helpful for training liu2024decade . It is worth mentioning that training inefficiency and instability are significant challenges when training LLMs, which is both computational expensive and time-consuming. On the other hand, compared with previous RL scenarios in robot control and video games openai2016gym , LLM setting poses new characteristic: fast simulation, deterministic environment, and trajectory-level reward li2024remax . In this scenario, sampling desired trajectories from the environment is relatively more efficient.
To this end, we propose our new RL algorithm, Minimalist Policy Gradient Optimization (MyGO), which shifts the burden from model training to trajectory sampling, in which we can fully leverage the fast and relatively inexpensive simulation of LLMs kwon2023efficient ; zheng2024sglang . Unlike previous methods, MyGO disentangles the RL training process into two phases: sampling and training. During sampling, MyGO directly samples trajectories from the asymptotically optimal distribution and consequently allows the use of Maximum Likelihood Estimation (MLE) to optimize the policy function. Notably, optimizing a model via MLE is simple, stable, and well established as a regular training technique widley used in unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training .
Sampling
We first characterize the target optimal distribution $\pi^{*}$. Alternative to the original objective $\mathcal{J}$, we consider the expected reward with entropy regularization $\mathcal{J}^{\prime}=\mathbb{E}[r(\tau)]+\alpha\mathbf{H}(\pi)$ with $\alpha\in\mathbb{R}^{+}$, which, as a Boltzmann distribution ziebart2010modeling , has a closed-form for the corresponding optimal policy, denoted as ${\pi^{\prime}}^{*}$:
[…]
where $\alpha$ is the temperature and $Z(\alpha)$ is the partition function that normalizes the distribution over the space of all possible trajectories $\mathcal{T}$. With a sufficiently small $\alpha\to 0$, we have $\mathcal{J}^{\prime}\to\mathcal{J}$, thus ${\pi^{\prime}}^{*}\to\pi^{*}$. As $\alpha\to 0$, ${\pi^{\prime}}^{*}$ becomes sharp and highly skewed, concentrating its probability mass on trajectories yielding high rewards. Therefore, given a sufficiently large threshold $K$ close to $\sup_{\tau\in\mathcal{T}}r(\tau)$, sampling trajectories $\tau$ with $r(\tau)>K$ is asymptotically equivalent to sampling from the ${\pi^{\prime}}^{*}$ and thus effectively approximating to sampling from the optimal policy ${\pi}^{*}$ of $\mathcal{J}$.
In practice, we relax the choice of $K$ by setting an update role of $K:=\max(K^{\prime},(\bar{r}_{n}/r_{\text{sup}}+1)\cdot r_{\text{sup}})$, where $r_{\text{sup}}$ denotes $\sup_{\tau\in\mathcal{T}}r(\tau)$, $\bar{r}_{n}$ is the empirical average reward of a batch of $n$ sampled trajectories, and $K^{\prime}$ is the previous threshold value which is initialized as a moderate value depending on the choice of dataset. As the policy $\pi_{\theta}$ improved, the $K$ will be asymptotically close to $\sup_{\tau\in\mathcal{T}}r(\tau)$, thus making the sampling distribution close to ${\pi^{\prime}}^{*}$.
We highlight that this approach is practically feasible for the reward landscapes for tasks of LLM agent, which differ significantly from that in complex robotic control problems. For LLM agents, trajectories frequently culminate in outcomes that are clearly classifiable as successful or unsuccessful. Empirically, we observe that well-prompted LLMs can readily generate a significant proportion of such successful trajectories, which makes directly sampling from $\pi^{*}$ feasible.
Let $\pi^{>K}$ denote the distribution that consists of trajectories with $r>K$. To clarify, we provide two propositions as follows, where the proofs can be found in the Appendix:
Proposition 3.1.
Given $\delta\in\mathbb{R}$, $\alpha\in\mathbb{R}$, there exists $K\in\mathbb{R}^{+}$ such that the following inequality holds:
$\mathds{D}_{\text{KL}}(\pi^{>K}||\pi^{*})<\delta\;,$
Proposition 3.2.
The following property holds when $\alpha\to 0$:
[…]
Proposition 3.1 states that if $K$ is sufficiently large, that the $\pi^{>K}$ can always approximate $\pi^{*}$. Proposition 3.2 says the variance of $\pi^{>K}$ is small, considering that $Z$ is usually a large value. These two propositions guarantee the proximity and stability of our proposed sampling strategy.
Learning
Once a sufficiently large dataset of optimal trajectories $\mathcal{D}$ has been collected, the optimization of policy $\pi_{\theta}$ can be simplified to maximizing the log-likelihood $\mathcal{L}$ of these trajectories. This yields the following MLE gradient:
$\nabla_{\theta}\mathcal{L}$ $=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\nabla_{\theta}\log\pi_{\theta}(\tau_{i})=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\sum_{j=0}^{|\tau_{i}|}\nabla_{\theta}\mathds{1}(o_{j}^{i})\log\pi_{\theta}(o_{j}^{i}|o_{<j}^{i})\;,$ (6)
where $o_{j}^{i}$ denotes the $j$-th token from $\tau_{i}$ and $o_{<j}^{i}$ denotes the sequence of tokens preceding $o_{j}$, $\mathds{1}(o_{j})$ is a characteristic function indicating whether $o_{j}$ is a token from the agent. The last equality holds because the action likelihood in language setting could be further elaborated with token likelihood, i.e, $\log\pi_{\theta}(a_{t}|s_{t})=\sum_{j}\log\pi_{\theta}(o_{j}|o_{<j})$.
This MLE objective, Eq. 6, is equivalent to a standard cross-entropy loss function, widely used in both unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training of language models. Compared to PPO-style algorithms, Eq. 4, MLE does not incorporate an advantage function, policy ratio clipping, and importance sampling. Similarly, unlike methods such as VPO, Eq. 3, MLE directly uses the curated data $\mathcal{D}$ for training and does not explicitly use reward values in the loss computation. We empirically observed that MLE leads to more stable training. We explain this increased stability by several factors: (i) the dataset $\mathcal{D}$ consists of trajectories sampled from a more stationary (approximated optimal) distribution, reducing variance compared to on-policy RL where the data distribution shifts rapidly; (ii) the direct use of a log-likelihood objective (cross-entropy) is often well-behaved and less sensitive to reward scaling issues that can affect RL algorithms; and (iii) the avoidance of importance sampling ratios, which can introduce instability if they become too large or small. A detailed theoretical justification for these empirical observations is left for future work.
Compared with previous works, like RAFT dongraft and STaR zelikman2022star , who try to simplify RL in LLM post-training, we use a progressive threshold to select data. More importantly, we provide a theoretical justification that elucidates why this approach, yet relatively simple, yields strong performance, offering novel insights into its efficacy.
4 Experiments
4.1 Environment Design
Datasets and Evaluation Metrics
To evaluate the effectiveness of our proposed agent workflow (Mujica) and training method (MyGO) on complex question answering tasks, we utilize four representative benchmark datasets: 2Wiki-MultihopQA (2Wiki) (ho2020constructing, ), QALD-10 (QALD-10, ), HotpotQA (yang2018hotpotqa, ), and MuSiQue (trivedi2021musique, ). 2Wiki and QALD-10 are KBQA benchmarks; HotpotQA and MuSiQue are text-based QA dataset. We use 2Wiki and HotpotQA for training, QALD-10 and MuSiQue for zero-shot evaluation. We assess QA performance using Exact Match (EM) and F1 score. For each dataset, these metrics are computed using the official evaluation scripts provided by the respective dataset creators. Comprehensive dataset statistics and precise definitions of the evaluation metrics are detailed in Appendix.
Environment Setting
To evaluate our agent’s capabilities across different data modalities, we derive two experimental environments from the 2Wik dataset: 2Wiki-KG and 2Wiki-Text. This setup allows us to agent’s performance on both structured (KGs) and unstructured (text) data. We use the F1 score as reward function for all environments.
The 2Wiki-KG environment is centered around the Wikidata KG (Wikidata, ). A KG is composed of factual statements represented as triples, such as (Donald Trump, position held, President of the United States). When the agent visits an entity, it can access all triples for which that entity is the head. The agent’s task involves traversing the KG by interpreting the semantic relationships encoded in these triples. For example, to answer the question "Who is the mother of the director of film Polish-Russian War (Film)?", the agent must identify and follow the correct reasoning path, such as: Polish-Russian War, director, Xawery Żuławski, mother, Małgorzata Braunek. To emphasize the reasoning aspect, we follow prior work (ToG, ; ToG2, ) by providing the agent with the gold topic entities extracted from the dataset as the entry.
We establish the 2Wiki-Text environment to further evaluate our agent’s ability to interact with text-based search engines. The original 2Wiki dataset provides ten supporting passages for each question, which contain the necessary information to derive the answer and the "distractor setting" is designed to test multi-hop reasoning abilities without retrieval. In our 2Wiki-Text environment, we adapt this by aggregating all passages to form a large corpus. This approach, following prior work (IRCoT, ; aguerrebere2023similarity, ), simulates an open-domain retrieval task. The agent’s objective is to answer questions by issuing search queries and then synthesizing information from the retrieved passages.
Beyond the 2Wiki environments, we extend our text-based evaluation using the HotpotQA dataset, from which we construct two additional settings: Hotpot and Hotpot-Kimi. In Hotpot, we folow the vanilla setting that uses the Wikipedia dump provided by the dataset yang2018hotpotqa , which only includes the title and the introductory paragraph. However, we found that the retriever karpukhin2020dense is the bottleneck for QA performance in the HotpotQA dataset. To isolate and more directly assess the agent’s reasoning capabilities independent of retrieval imperfections, we introduce the Hotpot-Kimi environment. Similar to but different from the distractor settings (yang2018hotpotqa, ; ho2020constructing, ), in the Hotpot-Kimi environment,the agent is directly provided with the ten gold supporting passages associated with each question, thereby guaranteeing access to all necessary evidence. However, its primary purpose here is to establish an idealized retrieval condition. Therefore, the gold passages are not directly fed into the QA-agent like distractor setting. Detailed examples and prompts for each environment are provided in Appendix.
4.2 Main Results
Table 1: Comparative Performance of Models on 2Wiki-KG and 2Wiki-Text Environments.
Type Model 2Wiki-KG 2Wiki-Text
EM F1 EM F1
Few-Shot GPT-4.1 78.50 84.24 30.70 52.14
Few-Shot Qwen2.5-7B 57.49 62.31 12.41 19.71
WarmUp 74.93 80.74 50.18 56.25
MyGO 77.63 84.15 53.17 59.62
Few-Shot Llama3.1-8B 77.70 82.09 37.83 46.53
WarmUp 82.00 85.90 58.03 64.81
MyGO 85.93 91.61 58.88 65.88
We evaluate our QA agent under three distinct settings: Few-shot, WarmUp, and MyGO. In the Few-shot setting, the agent is provided with a small number of illustrative examples via in-context learning within the prompt. Following common practice, we then perform a WarmUp phase. This phase is to adapt the base model beyond reliance on few-shot exemplars. For this purpose, we fine-tune the model on 1k samples with EM of 1. The MyGO training then commences, using this warmed-up model as its initialization. We experiment with two open-source models: Qwen2.5-7B-Instruct (Qwen) and Llama-3.1-8B-Instruct (Llama).
Table 1 presents the performance of these models on the 2Wiki dataset. Agents trained with MyGO consistently demonstrate strong performance. In the 2Wiki-KG environment, MyGO-Qwen achieves performance on par with a few-shot prompted GPT-4.1, while the Llama model significantly surpasses this baseline. In the 2Wiki-Text environment, both MyGO-trained Qwen and Llama models substantially outperform the GPT-4.1 few-shot baseline. Despite the same underlying questions, overall performance in the 2Wiki-Text environment is notably lower than in 2Wiki-KG. This disparity can be attributed to the differing data modalities and interaction paradigms. Interacting with KG may more closely resemble navigating discrete choices (e.g., selecting entities or relations), potentially reducing the burden on complex retrieval or synthesis from noisy search results. Conversely, text-based interaction requires robust open-retrieval over the corpus where explicit relationships between retrieved passages are often absent, demanding more sophisticated synthesis capabilities. Nevertheless, the consistent improvements observed affirm MyGO’s efficacy across these varied settings.
Table 2: Comparative Performance of Qwen on Hotpot and Hotpot-Kimi Environment.
Type Model Hotpot Hotpot-Kimi
EM F1 EM F1
Few-Shot GPT-4.1 29.80 48.09 33.50 53.35
Few-Shot Qwen2.5-7B 15.77 27.96 15.60 32.70
WarmUp 40.55 52.35 52.51 66.04
MyGO 41.54 53.79 54.07 68.48
Table 3 details the results on the HotpotQA dataset. Notably, even the WarmUp models exhibit strong performance, outperforming the few-shot GPT-4.1 baseline. However, the performance gain from MyGO over the WarmUp models is less pronounced in the Hotpot environment. We attribute this observation to two limitations inherent in this setup. Firstly, the efficacy of the dense retriever can be a bottleneck; essential evidence may not be surfaced for some queries. Secondly, the HotpotQA evaluation protocol has no answer aliases. This is problematic because LLMs frequently generate answers that are semantically equivalent to the gold answer but lexically different, which are penalized by strict string-matching metrics. To mitigate the impact of imperfect retrieval, we further evaluate using the Hotpot-Kimi environment, which is designed to simulate near-perfect retrieval by providing gold supporting passages. In this idealized Hotpot-Kimi setting, MyGO demonstrates substantial and consistent performance improvements over the WarmUp models, underscoring its effectiveness when reasoning capabilities can be more directly assessed. A more comprehensive comparison is provided in Appendix.
4.3 Out-Of-Distribution Evaluation
[…]
Type Model H $\to$ M 2W $\to$ Q
EM F1 EM F1
Few-Shot Qwen 2.5-7B 5.34 17.81 27.92 37.11
MyGO 26.11 35.92 39.85 49.73
To assess generalization capabilities, we conduct out-of-distribution (OOD) evaluations where models are RL post-trained on one dataset and evaluated on another one. Specifically, we examine cross-dataset transfer from HotpotQA to Musique (text-based), and from 2Wiki-KG to Qald-10 (KG-based). As shown in Table 3, MyGO demonstrates strong OOD performance. As presented in Table 3, models trained with MyGO demonstrate robust OOD generalization. Specifically, applying MyGO on the source datasets leads to significant performance improvements on the respective target OOD datasets when compared to the few-shot baseline. These findings indicate that MyGO does not overfit to the training data; rather, it appears to foster the acquisition of more generalizable reasoning and decision-making strategies. This capacity for effective generalization to unseen datasets aligns with a primary motivation for employing RL in the development of intelligent agents.
5 Related Works
MHQA with textual corpus.
Retrieval-based methods dominate the landscape of textual MHQA since supporting contexts from external corpora can effectively compensate for deficiencies in the inherent knowledge of language models and mitigate hallucinations. Traditional approaches primarily utilized sentence transformers bgem3 for dense retrieval or BM25 BM25 for keyword matching. Recently, extensive research has focused on optimizing the QA pipeline, particularly through the decomposition of complex questions IM-RAG ; ReAct , query rewriting and refinement RQ-RAG ; RaFe , and the iterative retrieval of relevant evidences in multiple steps SelfRAG ; ReSP ; Iter-RetRAG ; IRCoT .
Knowledge-based question answering.
Early approaches treat KBQA as a semantic parsing task, which generates and executes SPARQL queries lan-jiang-2020-query ; SPARQL2021 ; KD-CoT . These methods usually suffer from syntax errors and inaccurate entity labels, resulting in limited accuracy. Recently, researchers have shifted towards leveraging RAG and the advanced semantic understanding capabilities of LLMs to generate answers RRA . LLM-based KBQA approaches can be generally categorized into two types: iterative exploration and subgraph reasoning. Iterative exploration methods ToG ; ToG2 ; ToG-I ; KG-CoT ; PoG ; FiDeLiS ; GoG ; KARPA ; Paths-over-Graph ; BeamRetrieval ; DoG progressively retrieve and verify KG elements through multi-step reasoning, balancing interpretability with computational overhead. These approaches rely heavily on LLMs’ inherent reasoning capabilities due to the lack of supervised intermediate steps in most datasets. In contrast, Subgraph reasoning methods RoG ; GNN-RAG ; SubgraphRAG ; READS ; GCR extract relevant KG substructures in a single retrieval step for LLM-based answering, enabling efficient fine-tuning but suffering from the incompleteness of retrieved subgraphs and underlying knowledge bases.
6 Conclusion
In this work, we proposed Mujica, an interactive conversational framework for multi-hop question answering, and introduced MyGO, a novel RL optimization method that overcomes the inefficiencies of Proximity Optimization-based approaches. Mujica leverages a planner-worker structure to efficiently decompose and resolve complex questions, while MyGO improves the training efficiency of proximity optimization methods by shifting the burden from clipping importance ratio and tweaking advantage functions to trajectory sampling. Our experiments demonstrate that the proposed framework not only achieves significant performance improvements across multiple QA datasets but also exhibits strong inductive reasoning capabilities, highlighting its scalability and potential to generalize effectively to real-world scenarios.
Appendix A Problem Definition
The MHQA task involves answering a complex question $q$ by performing multi-step reasoning over multiple pieces of supporting contexts $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{|\mathcal{C}|}\}$ to derive the final answer $A_{q}$. Based on the form of supporting contexts, MHQA can be further breakdown into two sub-tasks, text-based MHQA and Knowledge-based Question Answering (KBQA).
Text-based Multi-hop Question Answering
In text-based MHQA, the supporting contexts are paragraphs scatter in one single document or multiple passages. In practice, the text-based MHQA task has two different settings. The distractor setting provides a pre-extracted set of supporting paragraphs $\mathcal{T}_{q}$ for each question $q$, which includes both relevant information and distractors. In contrast, under the full-corpus setting, all questions share an open textual corpus $\mathcal{T}$ (e.g. Wikipedia). In this work, we focus on the full-corpus setting, since it is closer to the real-world scenario.
Knowledge-based Question Answering
Under the setting of KBQA, the supporting contexts are retrieved from a knowledge graph (KG). KGs consist a large set of entities $\mathcal{E}$ and their relations $\mathcal{R}$, typically represents as triples $\mathcal{T}$ in the form of $(e_{h},r,e_{t})$, where $e_{h}$ and $e_{t}$ denote the head and the tail entity, respectively, and $r$ is the relation connecting them. To answer a question $q$, the model needs to identify the topic entities mentioned in the question, and traverse the KG by chaining multiple triples together to obtain the final answer $A_{q}$.
Appendix B Mathematical Proofs
See 3.1
Proof.
[…]
where $\mathds{1}(r>K)$ is the characteristic function filtering out trajectories whose reward is less than $K$. The above statement tells that $\pi^{>K}$ is actually a conditional probability of $\pi^{*}$, i.e., $\pi^{>K}(\tau)=\pi(\tau|r(\tau)>K)$. Also, the above statement allows us to compute the KL divergence easily:
$\mathds{D}_{\text{KL}}(\pi^{>K}||\pi^{*})=\log\frac{Z}{Z^{>K}}\;.$ (8)
Notably,
[…]
where $M=\sup(r)$ and the last equality is obtained by the mean value theorem. In summary, if $K<M-\frac{Z}{\exp(\delta)\exp(r^{\prime}/\alpha)}$, then the proposition holds.
See 3.2
Proof.
First, we expand the variance by regarding $\pi^{>K}$ as a distribution with respect to $r$. Note that all $\mathrm{Var}$ and $\mathbb{E}$ here are on $\pi^{>K}$ and $M=\sup(r)$.
[…]
Note that we can expand the integral at $\alpha=0$ with a Taylor series:
$\mathrm{Var}[r]=$ (11)
$\frac{e^{(K+M)/\alpha}(-2\alpha^{2}+O(\alpha^{6}))+\left(e^{(2K)/\alpha}+e^{(2M/\alpha)}\right)\left(\alpha^{2}+O(\alpha^{6})\right)+(K-M)^{2}(-e^{(K+M)/\alpha})}{\left(e^{(K/\alpha)}-e^{(M/\alpha)}\right)^{2}}\;.$
Notice the fact that $Z^{>K}(\alpha)=\alpha(\exp(M/\alpha)-\exp(K/\alpha))$, the proposition holds. ∎
Appendix C Datasets and Metrics
To evaluate MyGO’s effectiveness on complex question answering, we employ four representative datasets: 2WikiMultihopQA (2Wiki) [19], QALD-10 [69], HotpotQA [80], and MuSiQue [67]. 2Wiki utilizes both structured (Wikidata) and unstructured (Wikipedia) data to create multi-hop questions. QALD-10 was initially designed for Multilingual Question Answering over KG but late was widely used as a benchmark on KBQA [63]. The other two datasets focus on multi-document and multi-hop reasoning. HotpotQA features complex questions requiring reasoning across multiple documents. MuSiQue consists of questions demanding two to four reasoning hops, constructed from five existing single-hop QA datasets.
Appendix D Implementation Details
D.1 Sampling Strategies
Reward Criterion
Ideally, we would use the “correctness” of predicted answers as the reward criterion. However, this “correctness” is not directly accessible in practice. The ground truths provided in QA task datasets are in fact the golden answers. The failure to achieve an exact match - a widely used metric in QA tasks - between the generated answer and the golden answer does not necessarily indicate the semantic “incorrectness”. For example, for a question in HotpotQA “John T. Mather Memorial Hospital is located in a village in what Suffolk County, New York town?”, the generated answer “Brookhaven” fails to achieve an exact match with the golden answer “Town of Brookhaven”, which would be considered correct by a human judge.
Therefore, we adopt the F1 score as our evaluation criterion and consider a generated answer correct if it achieves a sufficiently high F1 score close to 1.
Note that some works utilize more complicated non-standard criteria, e.g., LLM-as-Judge [10, 62], Cover Exact Match (CEM) [62, 60]. Since this is not the main focus of this work, we simply follow the standard criterion widely used in previous works.
Progressive Selection Threshold
As stated in section 3.3 as well as the aforementioned, we aim to directly sample trajectories $\tau$ for $r(\tau)>K$ with a sufficiently large $K$.
However, this might lead to very low sample efficiency when the current policy is not sufficiently strong or when the tasks are particularly challenging. As a compromise, we consider a progressive selection threshold defined as $K:=\max(K^{\prime},(\bar{r}_{n}/r_{\text{sup}}+1)\cdot r_{\text{sup}})$, where $r_{\text{sup}}$ denotes $\sup_{\tau\in\mathcal{T}}r(\tau)$ and $\bar{r}_{n}$ represents the empirical average reward of the trajectories for a batch of $n$ questions at the first pass. This approach balances both the difficulty of sampling satisfactory trajectories and the quality of the sampled trajectories. As the policy improves during the MyGO training procedure, $K$ will asymptotically grow to an acceptable value, thereby satisfying Proposition 3.1.
Multiple-Attempt Sampling
For a batch of $n$ questions, we would iteratively sample trajectories for them. For a question, the sampling procedure will stop if $m$ satisfying trajectories are sampled or reach the predetermined maximum sampling attempts. With vLLM engine [26], we could accelerate the sampling procedures in parallel efficiently with multi-threading/multi-processing.
D.2 Training
Online Off-policy Optimization
In MyGO, we primarily focus on online off-policy optimization. Given a total of $N$ data examples (i.e., questions), we divide them into $T$ batches for $T$ update iterations. For each iteration, we sample trajectories (including answers) from $N/T$ questions to serve as pseudo-labels and optimize the models via MLE through a standard SFT pipeline (e.g., LLaMA-Factory [90]). After completing one update iteration, we proceed to the next sampling iteration.
Notably, due to the off-policy strategy, we can parallelize the sampling and training processes after the first iteration, when a sufficient number of GPUs are available.
Trajectory Selection for Training
In the Mujica framework, the planner and workers collaborate to reach the final answer. This interaction generates a single planner trajectory and multiple worker trajectories. It is important to clarify that in Section 3.3, the term "trajectory" specifically refers to the planner’s trajectory. In contrast, the dialogue history for each worker is straightforward, comprising a single-turn interaction with the environment. The environment’s message to the worker consists of the question from the planner and passages from the retriever, while the worker’s message contains its response. If the reward associated with a planner trajectory exceeds a predefined threshold, that trajectory and its corresponding worker trajectories are selected as independent data for training.
Training Advances from Pre-training/SFT
Unlike other reinforcement learning techniques, MyGO’s training process can be framed as supervised fine-tuning (SFT) using self-generated data and labels. Consequently, we can directly leverage existing training advancements from pre-training and SFT techniques, including optimizers, learning rate schedulers, and hyperparameter settings.
D.3 Hyperparameters and Configurations
Sampling
For all environments and datasets, we sample a maximum of $m=3$ trajectories per question from a randomly selected subset of 20,000 questions from the training sets. The maximum number of sampling attempts is set to 16. During each iteration of the sampling process, we sample trajectories from 1,000 questions.
Warmup
To burn in the prompts, we introduce a warmup stage before performing MyGO, in which we use the few-shot model to randomly sample 300 trajectories with $\text{EM}=1$ from the training sets. We then fine-tune the pre-trained model with these trajectories following the same procedure as MyGO.
Configuration
Following the settings of regular SFT, we employ the AdamW optimizer [41] with a Cosine Annealing learning rate scheduler [40]. For the warmup phase, we configure the initial learning rate at 1e-5 and the minimum learning rate at 1e-6. In the MyGO training procedure, we set the initial learning rate to 5e-7 and the minimum learning rate to 1e-8. We leverage vLLM [26] as our inference engine and LLaMA-Factory [90] as our training framework.
Appendix E Comprehensive Comparison with Other Methods
Table 4: Evaluation Results on Three Representative Multi-hop Question Answering Datasets.
Method Backbone Retriever HotpotQA MuSiQue 2Wiki
EM F1 EM F1 EM F1
Direct Inference Qwen 2.5-7B - 18.3 - 3.1 - 25.0 -
CoT Qwen 2.5-7B - 9.2 - 2.2 - 11.1 -
RAG Qwen 2.5-7B E5 29.9 - 5.8 - 23.5 -
RAFT Llama 3.1-8B Dragon-Plus 41.0 51.6 13.8 24.0 39.4 45.8
RaFe GPT 4o-mini Google Search 40.6 55.4 11.8 23.8 36.2 39.3
HippoRAG GPT 3.5-turbo ColBERTv2 45.7 59.2 21.9 33.3 47.7 62.7
Iter-RetGen Qwen 2.5-7B FlashRAG 34.4 - 8.7 - 27.9 -
IRCoT Qwen 2.5-7B BM25 30.3 - 7.0 - 21.6 -
IterDRAG Gemini 1.5 Gecko 1B 38.4 49.8 22.6 35.0 44.3 54.6
Concurrent works
Search-o1 Qwen 2.5-7B E5 18.7 - 5.8 - 17.6 -
Search-o1 QwQ 32B Bing Web Search 45.2 57.3 16.6 28.2 58.0 71.4
Search-R1 Qwen 2.5-7B E5 37.0 - 14.6 - 41.4 -
RAG-Gym Llama 3.1-8B bge-base + BM25 44.1 56.8 - - 50.2 57.9
ReSearch Qwen 2.5-7B FlashRAG 43.5 - 22.3 - 47.6 -
Settings In-domain Out-of-domain In-domain
Mujica (few-shot) Qwen 2.5-7B bge-large-en-v1.5 15.77 27.96 5.34 17.81 12.41 19.71
Mujica w/ warm up Qwen 2.5-7B bge-large-en-v1.5 40.55 52.35 N/A N/A 50.18 56.25
Mujica-MyGO Qwen 2.5-7B bge-large-en-v1.5 41.54 53.79 26.11 35.92 53.17 59.62
In order to comprehensively evaluate the performance of our proposed solution, we compare Mujica-MyGO with a series of RAG systems. Apart from direct inference with pretrained LLM [78], we consider vanilla RAG [29], RAFT [87], RAFe [45], HippoRAG [17], Iter-RetGen [57], IRCoT [68], IterDRAG [83], as well as compatitative concurrent solutions Search-o1 [34], RAG-Gym [29] and ReSearch [10]. Considering the differences between the backbone LLMs and retrieval methods employed by each RAG system, it is inappropriate to compare the performance of different solutions solely based on their evaluation metrics.
Table 4 presents the experimental results on three representative multi-hop question answering datasets. We have the following observations:
Firstly, Mujica-MyGO is an effective and efficient multi-hop QA solution. Despite leveraging a 7B Qwen2.5 LLM as its backbone, Mujica-MyGO achieves state-of-the-art performance on the MuSiQue and 2Wiki datasets and delivers competitive results on HotpotQA. One may notice that some baseline methods outperform Mujica-MyGO on the HotpotQA dataset. However, these approaches typically rely on larger-scale backbone LLMs and/or more sophisticated retrieval methods. In general, larger backbone models (e.g., GPT and QwQ) tend to yield better results. In addition, incorporating advanced retrievers (e.g., HopRetriever [33], Dragon-plus [37]) or expanding the knowledge corpus via search engines (e.g., Google, Bing) can further enhance performance. Nevertheless, it is important to emphasize that improving retrieval is not the focus of this paper. In order to evaluate the effectiveness of the proposed method, we opt to adopt the official corpus released with the HotpotQA dataset [80].
Secondly, Mujica shows to be an exceptional multi-hop QA pipeline. The results of the Mujica w/ warm up variant demonstrate that even without relying on expensive reinforcement learning (RL), a small amount of fine-tuning data is sufficient for Mujica to outperform RAG baselines based on question decomposition and iterative retrieval. We attribute the advantages of our method to two key factors. On the one hand, the use of independent workers to process retrieved contexts effectively prevents irrelevant information from interfering with the planner’s reasoning process. On the other hand, formulating subquestions into a DAG fully accounts for the interdependencies between subquestions, ensuring reasoning coherence throughout the pipeline. However, every approach has its trade-offs. As a complex question-answering pipeline, Mujica imposes substantial demands on the LLM’s instruction-following and reasoning capabilities, making it difficult for small-scale pre-trained LLMs to achieve desirable results in few-shot settings.
Table 5: Comparison of Proximity Optimization-based Multi-hop QA Solutions.
Pipeline Post-training Method HotpotQA MuSiQue 2Wiki
Search-R1 PPO 37.0 14.6 41.4
RAG-Gym PPO 44.1 - 50.2
ReSearch GRPO 43.5 22.3 47.6
Mujica MyGO 41.54 26.11 53.17
Lastly, the experimental results reconfirm the effectiveness of MyGO. Compared with ReSearch (using GRPO), Search-R1 and RAG-Gym (using PPO), Mujica-MyGO achieves comparable or even better performance across three datasets with the commonly adopted bge-large-en-v1.5 dense retriever (see Table 5). This demonstrates that MyGO simplifies GRPO and PPO without compromising performance on multi-hop QA tasks. Moreover, MyGO does not require additional resources to hold a reference model or to compute advantages, striking a great balance between efficiency and performance.
Detailed comparisons between MyGO and other preference optimization algorithms (e.g., DPO, PPO, GRPO, etc.) across various tasks would be very interesting, however, this lies beyond the scope of this paper. We shall emphrasize that the focus of this paper is the multi-hop QA task. In our future works, we plan to extend our analysis to different reasoning tasks, including but not limited to mathematical problems, coding, and long-context reasoning.
Appendix F Further Study: Offline MyGO?
Designing an offline variant of MyGO is straightforward: we can use the initial policy after warmup to sample all required trajectories. In this work, we present a preliminary attempt using Qwen on the 2Wiki-KG and 2Wiki-Text datasets.
Figure 4: F1 score on 2Wiki-KG training
Figure 5: F1 score on 2Wiki-Text training
Table 6: Mujica+MyGO w/ Qwen Online v.s. Offline on 2Wiki-KG and 2Wiki-Text Environments.
Type Model 2Wiki-KG 2Wiki-Text
Mujica+ EM F1 EM F1
Few-Shot GPT-4.1 78.50 84.24 30.70 52.14
Few-Shot Qwen2.5-7B 57.49 62.31 12.41 19.71
WarmUp 74.93 80.74 50.18 56.25
MyGO-Offline 77.09 83.16 53.27 59.78
MyGO-Online 77.63 84.15 53.17 59.62
We note that our goal is not to compare online and offline RL, as this has been discussed in many prior works [65]. Instead, we just aim to present an alternative approach for implementing MyGO.
From Proposition 3.1, we can expect that the quality of samples obtained from MyGO-offline and MyGO-online should be comparable, as both methods sample directly from an asymptotically optimal policy. The former uses a fixed threshold $K$, while the latter can dynamically and progressively increase $K$, which aligns more closely with Proposition 3.1. In the demonstration experiments (as shown in Table 6), we observe that MyGO-Offline can also achieve outstanding performance, being slightly worse or comparable to MyGO-Online. This indirectly justifies the insight from Proposition 3.1.
Appendix G Extended Related Works
MHQA with textual corpus.
Retrieval-based methods dominate the landscape of textual MHQA since supporting contexts from external corpora can effectively compensate for deficiencies in the inherent knowledge of language models and mitigate hallucinations. Traditional approaches primarily utilized sentence transformers [8] for dense retrieval or BM25 [52] for keyword matching. Recently, extensive research has focused on optimizing the QA pipeline, particularly through the decomposition of complex questions [79, 81] and the iterative retrieval of relevant evidences in multiple steps [5, 22, 57, 68]. Most notably, ReAct [81] integrates retrieval actions into the reasoning trajectory exploration process, enabling dynamic interaction with external environments. IM-RAG [79] optimizes multi-round retrieval and reasoning via reinforced learning. Search-o1 [34] introduces an agentic RAG framework, empowering the language model to decide when to retrieve external knowledge and to refine retrieved contexts, achieving state-of-the-art performance.
Knowledge-based question answering.
Early approaches [27, 49, 71] formulates KBQA as a semantic parsing task, focusing on extracting entities and relations from natural language questions and translating them into SPARQL statements capable of executing logical queries over KGs [59, 2]. While these methods provide accurate and interpretable answers, they encountered significant limitations since SPARQL statements generated by language models often suffer from syntactic errors, including incorrect entity and relation labels as well as unexpected grammatical mistakes [82, 42, 30]. Moreover, the incomplete nature of KGs [32] can also lead to query failures, preventing these approaches from outputting correct answers.
Recently, researchers have shifted towards treating KBQA as an information retrieval (IR) task, leveraging RAG and the advanced semantic understanding capabilities of LLMs to generate answers based on factual knowledge stored in KGs [73]. LLM-based KBQA methods can be generally categorized into two types: iterative exploration and subgraph reasoning. Iterative exploration methods [63, 44, 86, 88, 9, 61, 77, 15, 64, 85, 30] adopt a beam search-like strategy to retrieve the most relevant relations and neighboring entities, followed by pruning operations guided by LLMs. These methods exhibit strong traceability, enabling responsible and interpretable reasoning, but often come with the drawback of high computational complexity [36]. Moreover, the absence of ground truth reasoning paths and intermediate nodes in most KBQA datasets limits the feasibility of applying SFT [42, 47]. As a result, their performance is inevitably constrained by the inherent reasoning capabilities of LLMs. In contrast, subgraph reasoning methods [42, 47, 31, 76, 43] retrieve all candidate reasoning paths related to the question from the local neighborhood of the topic entities and guide the LLM to generate answers based on these paths in a single step. This approach enable direct improvement of the LLM’s in-context learning capabilities through SFT. Nevertheless, its effectiveness is often constrained by the quality of retrieved context, which typically cover only a subset of the knowledge required to address complex questions, omitting other critical information [9]. Furthermore, due to the incompleteness of KGs, the existence of relevant reasoning paths within the subgraph is not guaranteed. In such scenarios, LLM may hallucinate and produce incorrect answers.
Simplifying Reinforcement Learning with LLMs
Recent research has focused on simplifying RL algorithms for LLMs. For instance, GRPO [58] simplifies PPO by replacing the value model with standard deviation. Similarly, ReMax [35] adapts VPO by using the logits of greedy decoding as a baseline for GAE. Other VPO modifications include RLOO [4], REINFORCE++[20], and GPG [12]. Another line of research replaces VPO with MLE through advanced sampling techniques. Within this category, RAFT [87] selects the highest-reward trajectory within a group, while RAFT++ [75] enhances this by incorporating rejection sampling. Both RAFT and RAFT++ can be viewed as specific instances of STaR [84]. MyGO aligns with this latter research direction. It selects trajectories using either a fixed or progressively improving threshold, a property ensuring that sampling approaches the optimal policy. We further provide a mathematical theory to explain the efficacy of this streamlined method. If we continue to simplify MyGO, it will converge to STaR [84]. However, our primary motivation stems from the goal of simplifying RL algorithms specifically for LLM agent training.
Appendix H Working Examples
Table 7: Prompt template of Planner for text-based multi-hop question answering
## System Prompts ##
*Your Task:** Answer the user’s question thoroughly and accurately.
Before answering, outline your plan and reasoning process within ‘<think>‘ tags. This includes understanding the question, identifying needed information, and planning your search queries.
You can use the search tool to get information.
Please put your search query in ‘<search>‘ tags, e.g., ‘<search> Who is the mother of Xawery Żuławski? </search>‘
Every ‘<search>‘ tag only accepts one search query.
But you can use multiple ‘<search>‘ tags if needed.
Note that your query should be explicit and specific.
It is difficult to search a query with reference.
For example, if you directly search ‘Who is the mother of the director of film Polish-Russian War‘, the search engine cannot find the information directly.
You should make a plan to decompose the question into multiple sub-queries.
Finally, integrate all the information and put your final answer within ‘<answer>’ tags.
## User ##
Question: Who succeeded the first President of Namibia?
## Assistant ##
<think>
The question asks about the person who succeeded the first President of Namibia.
1. I need to find who was the first President of Namibia.
2. I need to find who succeeded the first President of Namibia.
</think>
<search> Who was the first President of Namibia? </search>
## User ##
Obs: Sam Nujoma was the first President of Namibia.
## Assistant ##
<think>
Now I know that Sam Nujoma was the first President of Namibia.
I need to find who succeeded Sam Nujoma.
</think>
<search> Who succeeded Sam Nujoma? </search>
## User ##
Obs: Hifikepunye Pohamba succeeded Sam Nujoma as the President of Namibia.
## Assistant ##
<think>
The question asks about the person who succeeded the first President of Namibia.
Sam Nujoma was the first President of Namibia.
Hifikepunye Pohamba succeeded Sam Nujoma as the President of Namibia.
I have all the information to answer the question.
</think>
<answer> Hifikepunye Pohamba </answer>
Table 8: Prompt Template of Worker for Text-based Multi-hop Question Answering
## System Prompts ##
You will be given a question with a context.
The context includes multiple passages labeled by index [x].
Your task is to select the most relevant passage from the context based on the question.
Please state the reason why you choose the passage. Put your thinking process into the <think>…</think> tag.
Put the index of the selected passage into the <select>…</select> tag.
Based on the selected passage, you should answer the question. State the answer as a complete declarative sentence, mirroring the question structure.
Put your answer into the <sentence>…</sentence> tag.
If the passages have no information related to the question, put [-1] in the <select>…</select> tag
and try your best to answer the question according to your knowledge in the <sentence>…</sentence> tag.
## User ##
Question: What college did Kyeon Mi-ri attend?
Context:
$[0]$ Kyeon Mi-ri graduated from Seoul Traditional Arts High School in 1983, then studied Dance at Sejong University. She made her acting debut in 1984, and has since become active in television dramas, most notably as the arrogant and ambitious Lady Choi in the 2003 period drama "Dae Jang Geum" (or "Jewel in the Palace"), which was a hit not only in Korea but throughout Asia.
$[1]$ Shin Kyeong-nim was born on April 6, 1936 in North Chungcheong Province, South Korea. When he was young, Shin Kyong-rim frequented the people of Koreaś rural villages and collected the traditional songs they sang. Much of his poetry represents a modernization of things he heard then Shin Kyeong-nim graduated in English Literature from Dongguk University, from which time he strove to become a creative writer. In 1955 and 1956, he made his formal literary debut with the publication of poems "Day Moon" (Natdal), "Reeds" (Galdae) and "Statue of Stone" (Seoksang). He taught elementary school in his hometown for a period of time, before moving to Seoul to work as an editor for Hyundae munhak and Donghwa Publishers. But his strong desire to create poetry continued.
$[2]$ Paige Ackerson-Kiely received a BA in Asian Studies from the University of New Mexico in Albuquerque. Prior to this achievement, she attended Beloit College in Beloit, Wisconsin, Marmara University in Istanbul, and Birzeit University in Birzeit, Palestine.
$[3]$ Han Seung-yeon was born on July 24, 1988, in Seoul, South Korea. She made her acting debut as a child actress in a bit part in "Dear Ends" (1993), "Summer Showers" (1995) and "Star in My Heart" (1997). She later left South Korea to study at Tenafly High School in New Jersey, United States. However, she withdrew from high school mid-course in order to pursue a singing career. After returning to South Korea, she debuted as a member of girl group Kara on March 29, 2007. During her time with the group, she passed a high school qualification exam, the College Scholastic Ability Test, and was accepted by Kyung Hee University, majoring in theater and film.
$[4]$ Myo Min Zaw studied English at the University of Yangon, where he became active in the pro-democracy group All Burma Federation of Student Unions (ABFSU). In December 1996, he participated in a student protest, and following the closing of Burma’s universities, remaining involved in the pro-democracy movement.
## Assistant ##
<think> The question asks about what college Kyeon Mi-ri attended.
Passage [0] clearly states that "Kyeon Mi-ri graduated from Seoul Traditional Arts High School in 1983, then studied Dance at Sejong University."
So I select passage [0]. </think>
<select> [0] </select>
<sentence> Kyeon Mi-ri attended Sejong University. </sentence>
Table 9: Prompt Template of Planner for Knowledge-based Question Answering
## System Prompts ##
Your task is answering questions. You will be given a question and some candidate entities. You should first make a plan about how to solve this question.
You can use the Search(entity, question) function to get information.
Please think thoroughly and label the final answer with <answer> *** </answer>.‘
## User ##
Question: Which film came out first, Blind Shaft or The Mask Of Fu Manchu?
Candidate: [0] Blind Shaft [1] The Mask Of Fu Manchu
## Assistant ##
<think>
To solve this problem, I need to:
1. Figure out when Blind Shaft came out.
2. Figure out when The Mask Of Fu Manchu came out.
3. Compare their dates.
I need to search information for both of Blind Shaft and The Mask Of Fu Manchu.
</think>
<action>
Search([0], "When did Blind Shaft come out?")
Search([1], "When did The Mask Of Fu Manchu come out?")
</action>
## User ##
Obs: Blind Shaft came out on 2003 and The Mask Of Fu Manchu came out on 1932.
## Assistant ##
<think> Ok. Right now I need compare their released date. 1932 is much earlier than 2003. Therefore, The Mask Of Fu Manchu came out first. </think>
<answer> The Mask Of Fu Manchu </answer>
Table 10: Prompt Template of Worker for Knowledge-based Question Answering
## System Prompts ##
Task: Extract relevant information from provided materials and present it in a specified format.
Instructions:
Read Carefully:
First, analyze the question to identify the key entity and requested information (e.g., "mother," "birth date").
Review all materials systematically.
Extraction Rules:
Use only direct matches from materials. Never paraphrase or infer relationships.
If multiple materials contain the answer, select all of them.
If no material matches, respond with “No relevant information found.”
Answer Format:
State the answer as a complete declarative sentence, mirroring the question structure.
Add the source ID in brackets at the beginning of the answer.
Use template:
<think> … let’s think step by step … </think>
<select> [X][Y] </select>
<sentence> Full sentence answer </sentence>
## User ##
Question: Who is the mother of Xawery Żuławski?
Materials:
[0] Xawery Żuławski, mother, Małgorzata Braunek
$[1]$ Xawery Żuławski, father, Andrzej Żuławski
$[2]$ Xawery Żuławski, family, Q63532193
$[3]$ Xawery Żuławski, family name, Q56541485
$[4]$ Xawery Żuławski, spouse, Maria Strzelecka
$[5]$ Xawery Żuławski, date of birth, 1971-12-22T00:00:00Z
$[6]$ Xawery Żuławski, sibling, Vincent Zulawski
$[7]$ Xawery Żuławski, place of birth, Warsaw
$[8]$ Andrzej Żuławski, child, Xawery Żuławski
$[9]$ Małgorzata Braunek, child, Xawery Żuławski
## Assistant ##
<think> The question asks me to find the mother of Xawery Żuławski. [0] says Xawery Żuławski’s mother Małgorzata Braunek, which excatly meets our need. </think>
<select> [0] </select>
<sentence> The mother of Xawery Żuławski is Małgorzata Braunek. </sentence>